OpenStudy (anonymous):

Pls Help!! Use the formula B=P(1+(r/n))nt which is used to calculate the final balance of an investment or a loan after being compounded. Following is a list of what each variable represents: B = the final balance P = the principal r = the annual percentage rate t = the time in years in which the principal is compounded n = the number of times the principal is compounded in a year Raul invests $40,000 in an account that is compounded annually. If, after two years, his balance is $45,796, what was the interest rate of the account?

OpenStudy (anonymous):

Are u helping me ...thank you , thank you , thank you

OpenStudy (anonymous):

yeah - sorry popcorn - i'll be back in about 5 minutes

OpenStudy (anonymous):

hi - you formula is incorrect. It should be B=P(1+(r/n))^nt - nt is an exponent so: 45796 = 40000(1 + r)^2 n=1 once a year (1+r)^2 = 45796 / 40000 = 1.1449 taking logs 2 ln(1+r) = ln 1.1449 ln (1+r) = ln 1.1449 / 2 1+r = 1.07 r = .07 = 7 % annual rate

OpenStudy (anonymous):

ok working thru it

OpenStudy (anonymous):

what is taking logs

OpenStudy (anonymous):

it means using the logarithm of the expression - its useful if the x is in the power live 2^x. so if got 2^x = 3.2 we can write ln 2^x = ln 3.2 so x ln2 = ln 3.2 and then x =ln3.2/ln2. hey - i could have done this another way - i could have expanded (1+r)^2 and solved for x. - but the log way is ok

OpenStudy (anonymous):

power like

OpenStudy (anonymous):

k i think my level is at the expansion level not log...would you be able to show me the other way??

OpenStudy (anonymous):

ok (1+r)^2 = 1.1.449 - hey - i've found another way we can take the square root of both sides 1 + r = sq rt 1.1449 =1.07 r = 1.07 - 1 = .07 I hope I haven't confused you with all these alternatives. I can't understand why I didn't see this easy way before!!

OpenStudy (anonymous):

got it thanks...if you are bored I have two other questions out there ;-)

OpenStudy (anonymous):

sorry popcorn its late here in uk and i've got an early start in the morning we are travelling to London to see my soccer team (Swansea) play a big game. I'll be here tomorrow evening.

OpenStudy (anonymous):

works for me...have fun

OpenStudy (anonymous):

bye